OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

need help anybody

OpenStudy (campbell_st):

you need to solve these equations simultaneously To eliminate y, multiply the 1st equation by 4 and the 2nd equation by -3 and then add the equations 8x - 12y = 12 -15x + 12y = -12 ------------------- -7x + 0 = 0 so the value of x = 0 the value of y would be y = -1 hope this helps

OpenStudy (anonymous):

thanks dude